OpenStudy (anonymous):

by substitution -x/6 + y/2=1/2 x/3-y/6=-3/4

OpenStudy (hba):

just get the value of x from one and put into another u will get y

OpenStudy (anonymous):

I think it becomes simple if we multiply 12 with both the equations first and then evaluate x or y for substitution..

OpenStudy (anonymous):

can u show me how waterineyes?

OpenStudy (anonymous):

Ditto to what @waterineyes said. Multiply EACH term by 12.

OpenStudy (anonymous):

Equations will become: -2x + 6y = 6 4x - 2y = -9 Now, evaluate x or y here..

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

Multiply EACH term in the top row that @waterineyes gave you by 2. Then add like terms to eliminate the x. -4x + 12y = 12 4x - 2y = -9 __________________ 0x + 10y = 3 So y = 3/10...plug that into ONE of the equations in order to solve for x.

OpenStudy (anonymous):

thanks @coreylvaughn

OpenStudy (anonymous):

You should be able to take it from there...

OpenStudy (anonymous):

@coreylvaughn we have to do it by substitution and not by elimination..

OpenStudy (anonymous):

No problem. Thank @waterineyes

OpenStudy (anonymous):

how to do it @waterineyes

OpenStudy (anonymous):

Ewwww...my apologies. Well at least now you know what y SHOULD come out to be once you've performed all substitutions.

OpenStudy (anonymous):

Pick one row. ONE row. And solve for ONE variable (i.e. get that variable ALONE)

OpenStudy (anonymous):

Plug whatever expression you get for that variable, into the SAME variable in the OTHER row. (Note: row = equation)

OpenStudy (anonymous):

Firstly divide by -2 in the first equation: we will get: x - 3y = -3 So, x = 3y - 3 Put it in second: 4(3y - 3) - 2y = -9 12y - 2y = -9 + 12 10y = 3 y = 3/10 Now for x: x = 3y - 3 x = (9/10) - 3 x = -21/10

OpenStudy (anonymous):

Thanks @waterineyes now I can solve some of the equation here

OpenStudy (anonymous):

Welcome dear...

OpenStudy (anonymous):

@waterineyes have a question in x dear?

OpenStudy (anonymous):

x=3y-3 right? and y= 3/10?

OpenStudy (anonymous):

I got it now. Thanks dear

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

got it dear... Thanks a lot!

OpenStudy (anonymous):

Sorry I was busy in something else.. You got y = 3/10 and you have x = 3y -3 to find x just put the value of y here..

OpenStudy (anonymous):

Welcome dear..